# A Survey on Monocular Re-Localization: From the Perspective of Scene Map Representation

### v1

@@ 4.4 Further Improvements
There are several other attempts to improve VL-MRL methods, except feature extraction and matching, and pose solver. Here we introduce some works with noticeable improvements.
@@
Cross Descriptor matching cross-desc : Traditional VL-MRL methods supposes local features used in mapping and localization stage are same, which blocks the usage of new local feature algorithm. For example, if we used SIFT sift in mapping while HardNet hardnet in localization, traditional VL-MRL solution fails. In cross-desc , the authors overcame this problem by proposing to translate different type of descriptors into other descriptors or joint embeddings so that different descriptors can be indirectly matched.
@@
Line Feature PL-Loc : Along with point-wise local features for matching, line-wise local features provide additional constraints in VL-MRL methods, especially in human-made environment. In PL-Loc , authors first presented a novel line segment descriptor and suggested the line feature can be complementary constraints to traditional point feature, as show in Fig. 11, thereby boosting localization performance on extreme scenarios where point features are biased or sparse.
Figure 11: Points and lines are complementary for better localization PL-Loc .
Dense CNN Matching InLoc ; PoseVerification ; Stumberg2019GNNetTG : Indoor environments are full of large textureless areas, where sparse local feature extraction methods detect very few features and extracted local features are clustered in small, well-textured parts in the image, leading to potentially unstable configurations for VL-MRL. To overcome this problem, InLoc InLoc and its extension PoseVerification use multi-scale dense CNN features for both image description and feature matching. GN-Net Stumberg2019GNNetTG is optimized with the proposed GN loss for training weather-invariant deep features, which is tailored for direct image alignment.
@@
VL-MRL without SfM meshloc : Traditional VL-MRL need to reconstruct scene map using SfM colmap in offline mapping stage, which is computational and memorial costly. MeshLoc meshloc explores a more flexible alternative based on dense 3D meshes that does not require local features matching between $\mathcal{I}_{r}$ to build the scene map. During localization stage, MeshLoc meshloc still perform local feature matching on 2D images and estimate pose by solving PnP prblem, but the only difference is $\mathcal{I}_{r}$ is online rendered by 3D mesh model.
@@
Map Squeeze SceneSqueezer : Storing pre-built scene map with visual landmarks can be prohibitively expensive for large-scale environments, especially on mobile devices with limited storage and communication bandwidth. In SceneSqueezer , authors designed a novel framework, SceneSqueezer, to compresses a scene while still maintaining localization accuracy. The scene $\mathcal{M}$ is compressed in three stages: 1) $\mathcal{I}_{r}$ for mapping are clustered; 2) a learned point selection module prunes the visual landmarks in each cluster taking into account the final pose estimation accuracy; 3) the descriptors of the selected visual landmarks are further compressed using learned quantization. By applying SceneSqueezer SceneSqueezer , the size of scene map is reduced to 31 MB compared to 7828 MB of HLoc hloc2 in Aachen Day-Night dataset benchmark6dof while achieving comparable localization performance.
@@
Pose Verification InLoc ; PoseVerification and Correction PoseCorrection : Some VL-MRL methods estimate a set of candidate poses for $\mathcal{I}_{q}$, a pose selection procedure need to be conducted. InLoc InLoc utilizes the 3D structure of indoor scene and explicitly renders a virtual view that shows how the scene looks like from the estimated query pose. Then, authors extracted RootSIFT RootSIFT on a regular grid or dense CNN features from synthesized view and measured the region similarity between $\mathcal{I}_{q}$ and synthesized image, which can be seen as a score of candidate poses. In PoseVerification , Taira et al. improved pose verification by integrating semantic constraint and trainable pose verification network. In PoseCorrection , to circumvent the sparsity of co-visible local features between $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$ and improve the accuracy, Hyeon et al. proposed pose correction to reorganize local features observed from the estimated pose so that more reliable candidate poses can be provided to pose selection.
@@
Open-sourced Toolbox hloc2 ; Kapture ; xrlocalization : To facilitate the development of VL-MRL researches, many great works have open-sourced the overall localization pipeline for the community, e.g., HLoc [^2] https://github.com/cvg/Hierarchical-Localization, Kapture-Localization [^3] https://github.com/naver/kapture-localization, and XRLocalization [^4] https://github.com/openxrlab/xrlocalization. By leverage these toolbox, researchers can quickly valid their new contributions or evaluate their new proposals without any unrelated efforts to build localization pipeline.
5 Point Cloud Map
@@
Visual landmark map in Sec. 4 has a large requirement of storage due to saved high-dimensional features descriptors. The sensitivity of appearance (e.g., illumination) also challenges mapping with monocular camera and affect the reconstruction of visual landmark map. As a alternate, LiDAR can provide direct and precise 3D perception of scene map geometry, and the point cloud map built by high-precision LiDAR always be illumination-invariant and commonly more accurate than visual map. And point cloud map only need to save 3D position of point cloud, alleviating the storage burden. The point cloud map can be represented as =$\mathcal{M}=\{\textbf{m}_{i}\}$ where $\textbf{m}_{i}$ is a 3D point in global frame $\mathcal{G}$, $\textbf{m}_{i}=<\textbf{p}^{\mathcal{G}}_{i},\textbf{l}_{i}>$, $\textbf{p}^{\mathcal{G}}_{i}$ is its 3D coordinates and $\textbf{l}_{i}$ is its intensity (intensity sometimes is missing). In practical usage, the 3D points are usually used in “clip” format, that is, given a virtual viewpoint $\textbf{x}^{\mathcal{G}}_{v}$, the 3D point observed in this view can be clustered in a set $\mathcal{P}^{\mathcal{G}}_{v}=\{\textbf{p}_{i},\textbf{l}_{i}\}$, so $\mathcal{M}$ is changed to =$\mathcal{M}=\{\textbf{m}_{i}\}$ where $\textbf{m}_{i}$ is a clip of 3D point and its associated viewpoint $\textbf{m}_{i}=<\mathcal{P}^{\mathcal{G}}_{v},\textbf{x}^{\mathcal{G}}_{v}>$. Since the wide deployment of LiDAR localization is hampered by the huge cost of high-precision LiDAR, localizing camera in LiDAR point cloud map (named as Image-to-Point cloud (I2P) localization for simplicity) is a newly emerged trend in MRL. However, the appearance information of camera and the geometry information of LiDAR is in different modality, the inherit difficulties in cross-modal data association challenges I2P Localization. In this section, we review cross-modal I2P localization using traditional geometry rules and using learning-based methods.
@@
Figure 12: The diagram of four typical cross-modal I2P localization methods. including geometry-based a) G-I2P mono_i2p_vo , and learning-based b) I2P-VPR, c) I2P-RPR, and d) I2P-MLoc.
5.1 Geometry-based Cross-modal Localization
@@
Geometry-based I2P (G-I2P) methods localize camera in LiDAR map by traditional geometry rules. In the G-I2P researches, Stereo camera is a natural pick because we can easily lift 2D image data to 3D by stereo matching, which makes the alignment between camera and point cloud map much easier stereo_i2p . Instead, Localizing monocular camera images in 3D point cloud maps is inherently more challenging, as they lack any depth or 3D information. For this, researchers turn to align $\mathcal{I}_{q}$ to $\mathcal{M}$ by lifting 2D visual measurements to 3D local construction mono_i2p_vo , digging geometry features mono_i2p_line ; Wolcott2014IROS , or modelling 3D structure in other way mono_i2p_SDF ; dsl ; Huang2020GMMLocSC . As an early and successful attempt, Caselitz et al. mono_i2p_vo proposed to associate the visual landmarks reconstructed from the monocular visual odometry orbslam with point cloud map, as show in Fig. 12 a). The 7 DoF transform was estimated in an ICP scheme. In Wolcott2014IROS , authors proposed to render synthetic images using intensity acquired from the LiDAR sensor so that pose estimation can be performed by maximizing the normalized mutual information between real $\mathcal{I}_{q}$ and synthetic images. Although identifying matched features across different modalities is proven challenging, the incorporation of line features has been demonstrated to aid in this objective in mono_i2p_line . Once the authors identified the co-visible lines in both 3D and 2D views, they designed a feature vector for each line, which incorporated various properties like length and orientation. Then, line feature matching between 2D and 3D views is performed to provide geometric constraint for pose estimation. Huang et al. modelled the dense structure as an Euclidean Signed Distance Field (ESDF) mono_i2p_SDF . The visual structure can then be aligned with the implicit surface. Ye et al. proposed DSLdsl where surfel constraints of point clouds are introduced into the direct photometric error. The monocular camera can be localized in a tightly-coupled BA framework with photometric error. More recently, GMMLoc Huang2020GMMLocSC uses a Gaussian Mixture Model (GMM) GMM to model the prior distribution of the point cloud map. Assisted by the GMM map, the performance of baseline orbslam2 in localization is significantly enhanced in challenging situations.
@@
Summary: Many attempts on G-I2P have been made and achieved impressive successes, but all of these works need to be integrated into a VO pipeline to obtain initial guess of pose, and they are hard to re-localize with only one monocular image. These methods tend to apply continuous re-localization in point cloud map for absolute pose estimation in order to add absolute constraint into VO system, alleviating the drift problem of long term exploration.
5.2 Learning-based Cross-modal Localization
@@
Cross-modal Visual Place Recognition (I2P-VPR): Recently, scholars have begun to explore deep learning-based methods to achieve I2P localization with only one monocular image. First, many researches adopt GF-VPR method (in Sec. 3.1) into this cross-modal I2P localization task and propose I2P-VPR algorithms. I2P-VPR methods are very similar to GF-VPR methods, but its reference “frames” $\mathcal{I}_{r(t)}$ is changed to local clips of point cloud map $\mathcal{P}^{\mathcal{G}}_{r(t)}$. For clearness, we also provide detailed pipeline of I2P-VPR methods here. Given the query image $\mathcal{I}_{q}$, a I2P-VPR algorithm can be written as a two-step pipeline, that is,
@@
Then, the pose of $\mathcal{I}_{q}$ can be approximated as the associated pose with $\hat{\textbf{m}}$.
For this problem, one straightforward solution is jointly training a 2D CNN $F_{\mbox{ImFeat}}(\cdot)$ for images and a 3D DNN $F_{\mbox{PcFeat}}(\cdot)$ for point clouds to create shared embeddings 2d3d_embed . However, this approach does not generalize well to unseen environments. Lukas2021ICRA_Spherical proposes to project images and point clouds into unit spheres and extract features through sphere CNN, which requires multiple images as input. Towards I2P localization robust to inconsistent environmental conditions, i3dLoc i3dloc matches equirectangular images to the 3D range projections by extracting cross-domain symmetric place descriptors. The 3D geometry features eliminates the condition-related features by a designed Generative Adversarial Network. Based on such features, i3dLoc i3dloc further design a spherical convolution network to learn viewpoint-invariant symmetric place descriptors. Similarly, AE-CrossModal focuses on correlating the information of 360-degree spherical images to point clouds. Attention mechanism is applied to let the network capture the salient feature for comparing images and point clouds. I2P-Rec Li2023I2PRecRI provides a new baseline for I2P-VPR that it leverages on depth estimation networks to recover point clouds from images so that cross-modal data is converted into the same modality, i.e., Bird-Eye-View (BEV) images. Using the BEV image as an intermediate representation, very simple global feature network (CNN encoder followd by NetVLAD layer) trained by small number of training data can achieve impressive localization performance.
Cross-modal Relative Pose Regression (I2P-RPR): After obtaining matched point cloud clip $\mathcal{P}^{\mathcal{G}}_{r}~{}\mbox{of}~{}\textbf{m}_{i}$ and its pose $\textbf{x}^{\mathcal{G}}_{r}$, many works further refine query pose by I2P-RPR that is similar to image-to-image RPR in Sec. 3.2, that is, estimate relative pose between $\mathcal{I}_{q}$ and $\mathcal{P}^{\mathcal{G}}_{r}$:
$\hat{\textbf{x}}^{r}_{q}=\mbox{RPR}(\mathcal{I}_{q},\mathcal{P}^{\mathcal{G}}_{r})$ (17)
so that the absolute pose of $\mathcal{I}_{q}$ can be calculated by:
@@
However, I2P-RPR methods are hard to achieve centimeter-level accuracy on large scale scene. So many proposals turn to re-localize camera based on 2D-3D matching between images and point cloud map, which is technically similar to VL-MRL methods in Sec. 4 and is also called image-to-point cloud registration.
@@
Cross-modal Matching-based Localization (I2P-MLoc): 2D3D-MatchNet 2d3dmatchnet is one of the earliest works focusing on image-to-point cloud registration for robot localization. The 2D and 3D key points are obtained by SIFT sift and ISS ISS respectively. Then, a neural network is introduced to learn the descriptors for key points. Finally, EPnP Lepetit2009EPnPAA is adopted to estimate the transformation between the image and the point cloud with the 2D-3D correspondences. Built upon CMRNet cmrnet , CMRNet++ cmrnet++ turn to regress displacements between images and projected depth of point cloud map instead of directly regressing pose. CMRNet++ cmrnet++ achieves improved performance with significant margin, indicating that matching-based methods should be a more reasonable choice than regression-based alternates for I2P localization. DeepI2P deepi2p splits the image-to-point cloud registration problem into a classification problem and an optimization problem. A cross-modality neural network is adopted to classify whether the points fall into the image frustum. The classification results are utilized to construct the cost function of the inverse camera projection. The optimization solver solves the transformation that minimizes the value of the cost function. CorrI2P CorrI2P designs a cross-modality network to extract the image-to-point cloud overlapping region and corresponding dense descriptors for the image and point cloud. CorrI2P CorrI2P constructs dense image-to-point cloud correspondences and uses iterative RANSAC-based EPnP Lepetit2009EPnPAA ; ransac to estimate the relative pose. EFGHNet EFGHNet adopts the divide-and-conquer strategy to divide the image-to-point cloud registration into four separate sub-networks. These sub-networks are responsible for the horizon and ground normal alignments, rotation estimation, and translation estimation, respectively. These mentioned above methods divide the image-to-point cloud registration into separate modules for robot localization. The separation makes the modules separately optimized and thus not able to refine the error of the previous modules. Thus, I2D-Loc I2DLoc apply BPnP module BPnP to calculate the gradients of the back-end PnP-based pose estimation process, enabling the model to be trained end-to-end with the supervision on estimated pose. More recently, I2PNet I2PNet utilizes a coarse-to-fine 2D-3D registration architecture to accurately regress pose. By combining both advantages of regression-based and matching based I2P localization methods, I2PNet I2PNet achieves centimeter-level localization accuracy.
@@
Some other works add visual features in point cloud map, which is quite similar to visual landmark map in Sec. 4 but the visual features of point cloud in map are online extracted during localization without offline storage. NRE introduces the Neural Re-projection Error (NRE) as a substitute for re-projection error in PnP problem. Given $\mathcal{I}_{q}$ and $\mathcal{P}^{\mathcal{G}}_{r}$ in map with $\mathcal{I}_{r}$, the method extracts dense CNN features from both images. The sparse descriptors of 3D point clouds are sampled from $\mathcal{I}_{r}$ based on image projection as Eqn. 12. For each 3D point, NRE computes dense loss maps and minimizes the loss with respect to the network parameters. The proposed NRE in NRE is differentiable not only w.r.t. the camera pose but also w.r.t. the descriptors. PixLoc pixloc is another work in this field. PixLoc pixloc also extracts multi-level features with pixel-wise confidences for $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$. The LM optimization then aligns corresponding features according to the 3D points in the map, guided by the confidence, from the coarse to the fine level. Compared with GN-Net Stumberg2019GNNetTG that also trained deep features via direct alignment, PixLoc pixloc leverages the power of differentiable programming, and the only used supervision in training is pose, without the requirement of accurate pixel-wise ground truth correspondences.
@@
Summary: The inherit difference between cross-modal data make I2P localization hard. Benefit from high-dimensional CNN features, matching across cross-modal data becomes possible. In general, I2P-VPR only serves to retrieve reference point cloud clip and provides an coarse pose estimation. I2P-RPR and I2P-MLoc could provide more precise pose estimation. I2P-RPR avoid explicit feature matching and pose solver, the training costs is cheap but it cannot achieve very high localization accuracy. Instead, I2P-MLoc performs best in cross-modal localization task, showing good prospects for development.
6 Vectorized High-Definition Map
In AD task, the compactness and accuracy are two vital characteristics of scene map. Therefore, the most widely used map in AD area has been HD Map since the its first appearance in late 2010. Commonly, the HD Map is created by mobile mapping system equipped with high-precision sensors including LiDAR, RTK and IMU at centimeter-level precision HDMap_review . The localization feature in the HD map includes dense point cloud and sparse map element. The dense point cloud in HD Map is similar to the raw point cloud map in Sec. 5. In this section, we focus on the sparse map elements usually represented as vectors with semantic labels. These elements correspond to road element or signatures in the real world, e.g., lighting poles, road markers, and lane lines. Such a semantic element-based map representation is much lighter than other scene map like point cloud map and visual landmark map while carries road elements with a high level of detail HDMap_review . MRL with HD map (HD-MRL) is concluded as an effective solution for mass-produced vehicles. The fundamental formulation of HD-MRL methods involve identifying semantic elements of HD map from images and then estimating pose by aligning the detected 2D element in the image with their corresponding 3D element in the HD map.

### v3

@@ IV-D Further Improvements
Except mentioned above methods, there are many attempts to improve VL-MRL methods in other perspectives. Here we introduce some works with noticeable improvements.
@@
Cross Descriptor matching [181]: Traditional VL-MRL methods supposes to use the same local features in mapping and localization stage, which blocks the development of new local feature algorithm. For example, if we used SIFT [56] in mapping while HardNet [122] in localization, traditional solution fails. In [181], the authors overcame this problem by proposing to translate different type of descriptors into other descriptors or joint embeddings so that different descriptors can be indirectly matched.
Dense CNN Matching [35, 182]: Indoor environments have large textureless areas where sparse local feature methods detect very few features and the extracted features are clustered in small, well-textured regions in the image. In this case, pose cannot be stably solved. InLoc [35] and [182] use MS dense CNN features for feature matching.
Figure 10: Points and lines are complementary for better localization [183].
@@
Line Feature [183]: Along with point-wise local features for matching, line-wise local features may futher help VL-MRL methods, especially in human-made environment. In [183], authors first presented a novel line segment descriptor and suggested the line feature can be complementary constraints to traditional point feature, as show in Fig. 10, thereby boosting localization performance on extreme scenarios where point features are biased or sparse.
@@
VL-MRL without SfM [184]: Traditional VL-MRL need to reconstruct visual landmark map using SfM [119] in offline mapping stage and the visual landmark map is generally 3D points, which is computational and memorial costly. MeshLoc [184] explores a more flexible alternative based on dense 3D meshes that does not require local features matching between $\mathcal{I}_{r}$ to build the scene map in mapping stage. During localization stage, it still perform local feature matching on 2D images and estimate pose by solving PnP problem, but the only difference is $\mathcal{I}_{r}$ is online rendered by 3D mesh model.
@@
Map Squeeze [185]: Storing pre-built visual landmark map can be prohibitively expensive for large-scale environments, especially on mobile devices with limited storage and communication bandwidth. In [185], authors designed a novel framework, SceneSqueezer, to compresses a scene while still maintaining localization accuracy. The scene $\mathcal{M}$ is compressed in three aspects: 1) $\mathcal{I}_{r}$ for mapping are clustered; 2) the visual landmarks in each cluster are pruned; 3) the descriptors of the selected visual landmarks are further compressed. By applying SceneSqueezer [185], the size of map is reduced from 7828 MB to 31 MB in Aachen Day-Night dataset [33] while achieving comparable localization performance.
@@
Pose Verification [35, 182] and Correction [186]: Some VL-MRL methods estimate a set of candidate poses for $\mathcal{I}_{q}$, a pose selection procedure needs to be conducted. Given an estimated query pose, InLoc [35] utilizes the 3D structure of indoor scene to explicitly renders a virtual image and measure its similarity to $\mathcal{I}_{q}$ as a score of query poses, which is called pose verification. In [182], Taira et al. improved the procedure by integrating semantic constraint and trainable pose verification network. In [186], Hyeon et al. proposed pose correction to reorganize local features observed from the estimated pose so that more reliable candidate poses can be provided to pose selection.
@@
Open-sourced Toolbox [24, 187, 188]: To facilitate the development of VL-MRL researches, many great works have open-sourced the overall localization pipeline for the community, e.g., HLoc [^2] https://github.com/cvg/Hierarchical-Localization, Kapture-Localization [^3] https://github.com/naver/kapture-localization, and XRLocalization [^4] https://github.com/openxrlab/xrlocalization. By leverage these toolboxes, researchers can quickly valid their new contributions or evaluate their new proposals without any extra efforts to build localization pipeline.
V Point Cloud Map
@@
Visual landmark map has a large requirement of storage for saving high-dimensional descriptors. The sensitivity of appearance also challenges mapping with monocular camera and affect the reconstruction accuracy of visual landmark map. Instead, LiDAR can provide direct and precise 3D perception of scene map geometry, and the point cloud map built by high-precision LiDAR always be illumination-invariant and commonly more accurate than visual map. And point cloud map only need to save 3D position of point cloud, alleviating the storage burden. The point cloud map can be represented as =$\mathcal{M}=\{\textbf{m}_{i}\}$ where $\textbf{m}_{i}$ is a 3D point in global frame $\mathcal{G}$, $\textbf{m}_{i}=<\textbf{p}^{\mathcal{G}}_{i},\textbf{l}_{i}>$, $\textbf{P}^{\mathcal{G}}_{i}$ is its 3D coordinates and $\textbf{l}_{i}$ is its intensity (intensity sometimes is missing). In practical usage, the 3D points are usually used in “clip” format, that is, given a virtual viewpoint $\textbf{x}^{\mathcal{G}}_{v}$, the 3D point observed in this view can be clustered in a set $\mathcal{P}^{\mathcal{G}}_{v}=\{\textbf{P}_{i},\textbf{l}_{i}\}$, so $\mathcal{M}$ is changed to =$\mathcal{M}=\{\textbf{m}_{i}\}$ where $\textbf{m}_{i}$ is a clip of 3D point and its associated viewpoint $\textbf{m}_{i}=<\mathcal{P}^{\mathcal{G}}_{v},\textbf{x}^{\mathcal{G}}_{v}>$. Since the wide deployment of LiDAR localization is hampered by the huge cost of high-precision LiDAR, localizing camera in LiDAR point cloud map (named as Image-to-Point cloud (I2P) localization for simplicity) is a newly emerged trend in MRL. However, the appearance information of camera and the geometry information of LiDAR is in different modality, the inherit difficulties in cross-modal data association challenges I2P Localization. In this section, we review cross-modal I2P localization using traditional geometry rules and using learning-based methods.
@@
Figure 11: The diagram of four typical cross-modal I2P localization methods. including geometry-based a) G-I2P [17], and learning-based b) I2P-VPR, c) I2P-RPR, and d) I2P-MLoc.
V-A Geometry-based Cross-modal Localization
Geometry-based I2P (G-I2P) methods localize camera in LiDAR map by traditional geometry rules. In the G-I2P researches, stereo camera is a natural pick because we can easily lift 2D image pixels to 3D point clouds by stereo matching, which makes the alignment between camera and point cloud map much easier [189]. Instead, localizing monocular camera images in 3D point cloud maps is inherently more challenging, as they lack any absolute depth or 3D information. For this, some methods align $\mathcal{I}_{q}$ to $\mathcal{M}$ by lifting 2D visual measurements to 3D local construction [17], digging geometry features [190, 191], or modelling 3D structure in other advanced way [192, 193, 194].
@@ V-A Geometry-based Cross-modal Localization
As an early and successful attempt, Caselitz et al. [17] proposed to reconstruct local visual landmarks by a monocular Visual Odometry (VO) [149] and simultaneously estimate initial pose. Using initial pose $\hat{\textbf{x}}^{\mathcal{G}}_{init}$, a local map is retrieved and cropped from global point cloud map. The resulting local reconstruction is aligned to local map in an ICP scheme so that a 7 DoF transform (including 6 DoF relative pose $\hat{\textbf{x}}^{init}_{v}$ with regard to local map and 1 DoF scale) can be estimated as shown in Fig. 11 a). MRL is achieved by $\hat{\textbf{x}}^{\mathcal{G}}_{v}=\hat{\textbf{x}}^{\mathcal{G}}_{init}*\hat{\textbf{x}}^{init}_{v}$. Then in [191], authors estimate pose by maximizing the normalized mutual information between real query image $\mathcal{I}_{q}$ and synthetic images rendered by LiDAR intensity. Although identifying matched features across different modalities is proven challenging, the incorporation of matched 2D-3D line-level features between query image and point cloud map has been demonstrated to aid in this objective in [190]. In DSL [193], Ye et al. introduced the surfel constraints of point clouds into the direct photometric error and estimated poses by a tightly-coupled BA framework. Huang et al. first modelled the dense structure as an Euclidean Signed Distance Field (ESDF) in [192] and then improved by modelling the prior distribution of the point cloud map by Gaussian Mixture Model (GMM) [195] in [194].
@@
Summary: Many G-I2P methods achieve considerable successes. With the absolute pose constraint by G-I2P and relative temporal constraint by VO, G-I2P methods can alleviate the accumulated drift problem of long-term VO and smoothen noisy estimation of absolute localization, achieving consistent and accurate pose estimation. But these works are limited to be coupled in a VO pipeline to obtain initial pose and hard to re-localize with only one monocular image. The monocular G-I2P solution is still an unsolvable problem.
V-B Learning-based Cross-modal Localization
@@
Cross-modal Visual Place Recognition (I2P-VPR): Recently, scholars have begun to explore DL-based methods to achieve I2P localization with only one monocular image. First, many researches adopt GF-VPR method into this cross-modal I2P localization task and propose I2P-VPR algorithms. I2P-VPR methods are very similar to GF-VPR methods, but its reference “frames” $\mathcal{I}_{r(t)}$ is changed to local clips of point cloud map $\mathcal{P}^{\mathcal{G}}_{r(t)}$. Given the query image $\mathcal{I}_{q}$, a I2P-VPR algorithm can be written as a two-step pipeline, that is,
@@
Then, the pose of $\mathcal{I}_{q}$ can be approximated as the associated pose of $\hat{\textbf{m}}$.
For this problem, one straightforward solution is jointly training a 2D CNN $\mbox{ImFeat}(\cdot)$ for images and a 3D DNN $\mbox{PcFeat}(\cdot)$ for point cloud clips to create matchable feature between cross-modal data [196]. However, this approach does not generalize well to unseen environments. [197] proposes to project images and point clouds into unit spheres and extract features through sphere CNN, which requires multiple images as input. Towards I2P localization robust to inconsistent environmental conditions, i3dLoc [198] matches equirectangular images to the 3D range projections of point cloud map. In i3dLoc [198], authors concludes visual condition-related information of scenes makes robust matching challenging, so we should eliminate condition-related features and extract condition-invariant 3D geometry features from visual inputs for matching to point cloud map only including geometry information. Therefore, a Generative Adversarial Network (GAN) is designed to extract features matchable to point cloud map. To improve performance of I2P-VPR, i3dLoc [198] also designs a spherical convolution network to learn viewpoint-invariant features. Similarly, [199] focuses on correlating the information of 360-degree spherical images to point clouds. Attention mechanism is applied to let the network capture the salient feature for comparing images and point clouds. I2P-Rec [200] provides a new baseline for I2P-VPR that it leverages on depth estimation networks to recover point clouds from images so that cross-modal data is converted into the same modality, i.e., Bird-Eye-View (BEV) images. Using the BEV image as an intermediate representation, very simple global feature network (CNN encoder followed by NetVLAD layer [19]) trained by few training data can achieve SOTA localization performance.
Cross-modal Relative Pose Regression (I2P-RPR): After obtaining matched point cloud clip $\mathcal{P}^{\mathcal{G}}_{r}~{}\mbox{of}~{}\textbf{m}_{i}$ and its pose $\textbf{x}^{\mathcal{G}}_{r}$, many works further refine query pose by I2P-RPR that is similar to image-to-image RPR in Sec. III-B, that is, estimate relative pose between $\mathcal{I}_{q}$ and $\mathcal{P}^{\mathcal{G}}_{r}$:
$\hat{\textbf{x}}^{r}_{q}=\mbox{I2P-RPR}(\mathcal{I}_{q},\mathcal{P}^{\mathcal{G}}_{r})$ (21)
so a fine absolute pose of $\mathcal{I}_{q}$ can be calculated by:
@@
However, I2P-VPR and I2P-RPR methods are hard to achieve centimeter-level accuracy on large scale scene. So many methods begin to explore 2D-3D matches between images and point cloud map, which is technically similar to VL-MRL methods.
@@
Cross-modal Matching-based Localization (I2P-MLoc): 2D3D-MatchNet [203] is one of the earliest works focusing on image-to-point cloud matching for robot localization. The 2D and 3D key points are obtained by SIFT [56] and ISS [204] respectively. Then, a neural network is introduced to learn the descriptors for cross-modal key points. Finally, EPnP [172] is adopted to estimate the pose of the image relative to the point cloud based on 2D-3D correspondences. Built upon CMRNet [18], CMRNet++ [205] turn to regress displacements between images and projected depth image of point cloud map instead of directly regressing pose. CMRNet++ [205] achieves improved performance with significant margin, indicating that matching-based methods should be a more reasonable choice than regression-based alternates for I2P localization. DeepI2P [206] splits the image-to-point cloud matching problem into a classification problem and an optimization problem. A cross-modality neural network is adopted to classify whether the points fall into the image frustum and the classification results are formatted as a cost function to measure how close the estimated pose to real one in optimization. CorrI2P [207] designs a cross-modality network to extract the image-to-point cloud overlapping region and corresponding dense descriptors for the image and point cloud. CorrI2P [207] constructs dense image-to-point cloud correspondences and uses iterative RANSAC-based EPnP [172, 118] to estimate the relative pose. EFGHNet [208] adopts the divide-and-conquer strategy to divide the image-to-point cloud matching problem into four separate sub-networks. These sub-networks are responsible for the horizon and ground normal alignments, rotation estimation, and translation estimation, respectively. These mentioned above methods divide the I2P-MLoc into two steps for robot localization, i.e., matching and localization. The separation makes the two steps separately optimized and thus not able to refine the error of the previous step. Thus, I2D-Loc [209] apply BPnP module [210] to calculate the gradients of the back-end PnP-based pose estimation process, enabling the model to be trained end-to-end with the supervision on estimated pose.
More recently, I2PNet [211] utilizes a coarse-to-fine architecture to accurately regress pose. By combining both advantages of I2P-RPR and I2P-MLoc methods, I2PNet [211] could achieve centimeter-level localization accuracy.
@@
Some other works add visual features in point cloud map, which is quite similar to visual landmark map in Sec. IV but the visual features of point cloud in map are online extracted during localization without the need of offline storage. [212] introduces the Neural Re-projection Error (NRE) as a substitute for re-projection error in PnP problem. Given $\mathcal{I}_{q}$ and $\mathcal{P}^{\mathcal{G}}_{r}$ in map with $\mathcal{I}_{r}$, the method extracts dense CNN features from both images. The sparse descriptors of 3D point clouds are sampled from $\mathcal{I}_{r}$ based on image projection as Eqn. 16. For each 3D point, [212] computes dense loss maps and minimizes the loss with respect to the network parameters. The proposed NRE in [212] is differentiable not only w.r.t. the camera poses but also w.r.t. the descriptors, so camera pose and descriptors can both be E2E optimized. PixLoc [213] is another work in this field. PixLoc [213] extracts multi-level features with pixel-wise confidences for $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$. The LM optimization aligns corresponding features according to the 3D points in the map, guided by the confidences, from the coarse to the fine level.
@@
Summary: The inherit difference between cross-modal data make I2P localization hard. Benefit from high-dimensional CNN features, matching across cross-modal data becomes possible. In general, I2P-VPR only serves to retrieve reference point cloud clip and provides an coarse pose estimation. I2P-RPR and I2P-MLoc could provide more precise pose estimation. I2P-RPR avoid explicit feature matching and pose solver, the training cost is cheap but its localization accuracy is limited. I2P-MLoc performs best in cross-modal localization task, showing good prospects for developments.
@@
VI Vectorized High-Definition Map
In AD task, the compactness and accuracy are two vital required characteristics of scene map. Therefore, the most widely used map in AD area has been HD Map since the its first appearance in late 2010. Commonly, the HD Map is created by mobile mapping system equipped with high-precision sensors including LiDAR, RTK and IMU at centimeter-level precision [214]. The localization feature in the HD map includes dense point cloud and sparse map element. The dense point cloud in HD Map is similar to the raw point cloud map in Sec. V. In this section, we focus on the sparse map elements usually represented as vectors with semantic labels. These elements correspond to road element or signatures in the real world, e.g., lighting poles, road markers, and lane lines. Such a semantic element-based map representation is much lighter than other scene map like point cloud map and visual landmark map while maintains highly detailed road elements useful to AD task [214]. MRL with HD map (HD-MRL) is concluded as an effective solution for mass-produced vehicles. The fundamental formulation of HD-MRL methods involve detecting semantic map elements from images and then estimating pose by aligning the detected 2D element with their corresponding 3D element in the HD map.